OpenStudy (anonymous):

How do you graph f(x)=(x-9)squared + 2

OpenStudy (anonymous):

you'll want to use the 'vertex form' of a parabola to help understand where to graph this at. does y=a(x-h)^2 + k look familiar?

OpenStudy (anonymous):

yes but im having trouble graphing it

OpenStudy (anonymous):

ok, so what do you get for the vertex of the parabola?

OpenStudy (anonymous):

I got 9,2 for my vertex

OpenStudy (anonymous):

great, so we can tell if the parabola is going to open up or down if the 'a' term is positive or neagative: if +a then opens up if -a then opens down

OpenStudy (anonymous):

I have that, but im not sure if my slope is right.. How do I find a?

OpenStudy (anonymous):

y =a(x-h)^2 + k f(x)=(x-9)^3 + 2 do you see where 'a' is for your equation? and what it is?

OpenStudy (anonymous):

No I dont know what a is.. is it x? therefore 1/1?

OpenStudy (anonymous):

f(x)=(x-9)^3 + 2 is the same as f(x)=1*(x-9)^3 + 2 so a = 1 and 'a' is positive, so the parabola opens up,

OpenStudy (anonymous):

we know the vertex and that the parabola opens up, the next thing to do is to input some numbers for x and see what y becomes. since we know the vertex is at (9,2) i'd us an x value of +1 and -1 from the x value 9, and see what y's I get, and plot those points

OpenStudy (anonymous):

Okay thanks I got the answer, I will give you medals

OpenStudy (anonymous):

thx